OpenStudy (realitywillslapyou):

I will give a medal! Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2. y = 2x - 8 y = 2x - 4 y = 2x + 7 y = 2x + 5

OpenStudy (phoenixfire):

What is the slope-intercept form?

OpenStudy (anonymous):

A Deff

OpenStudy (realitywillslapyou):

Thank you!

OpenStudy (anonymous):

no problem sir

OpenStudy (realitywillslapyou):

Sir?

OpenStudy (anonymous):

idk lol

OpenStudy (realitywillslapyou):

It's okay lol XD I'm just teasing you

OpenStudy (realitywillslapyou):

But it should by a "Ma'am" if anything

OpenStudy (anonymous):

oh haha it doesnt say gender i got confused there

OpenStudy (realitywillslapyou):

Yeah, I understand how about you just call me "dude"

OpenStudy (realitywillslapyou):

:D

OpenStudy (anonymous):

Miss will do :)

OpenStudy (realitywillslapyou):

Yeah.... I mean just look at my smiley face!(✿ノ◕ヮ◕)ノ*:・゚✧

OpenStudy (realitywillslapyou):

lol

OpenStudy (anonymous):

i like it lol

OpenStudy (realitywillslapyou):

Thanks!

OpenStudy (anonymous):

u on fb?

OpenStudy (realitywillslapyou):

No, don't have one

OpenStudy (anonymous):

oh ok. u should get one lol

OpenStudy (realitywillslapyou):

Why? I don't have any use for it.

OpenStudy (anonymous):

i only use it to contact ppl who i dont have numbers for or meet ppl

OpenStudy (realitywillslapyou):

Oh

OpenStudy (realitywillslapyou):

Can you help me with two more problems?

OpenStudy (realitywillslapyou):

Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 6 - x = 5x + 30 Equation #2 4x - 8 = 3x + 2 The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown.

OpenStudy (realitywillslapyou):

Which of the following is not equivalent to the formula for the circumference of a circle, C = 2 (pi)r ?

OpenStudy (anonymous):

1=A 2=D

OpenStudy (realitywillslapyou):

Thank you so very much!! :D

OpenStudy (anonymous):

your welcome!